Consider the case of Peaceful Book Binding Company

The CFO of Peaceful Book Binding Company is trying to determine the company’s WACC. He has determined that the company’s before-tax cost of debt is 9.60%. The company currently has $750,000 of debt, and the CFO believes that the book value of the company’s debt is a good approximation for the market value of the company’s debt.

The firm’s cost of preferred stock is 10.70%, and the book value of preferred stock is $45,000.
Its cost of equity is 13.50%, and the company currently has $500,000 of common equity on its balance sheet.
The CFO has estimated that the firm’s market value of preferred stock is $78,000, and the market value of its common equity is $880,000.

If PBBC is subject to a tax rate of 40%, Peaceful Book Binding Company’s WACC is   . (Note: Round your answer to two decimal places.)

Consider the case of Chilly Moose Fruit Producer

Chilly Moose Fruit Producer is considering a new project that will require an initial investment of $4 million. It has a target capital structure of 35% debt, 2% preferred stock, and 63% common equity. Chilly Moose Fruit has noncallable bonds outstanding that mature in 15 years with a face value of $1,000, an annual coupon rate of 11%, and a market price of $1,555.38. The yield on the company’s current bonds is a good approximation of the yield on any new bonds that it issues. The company can sell new shares of preferred stock that pay an annual dividend of $8 at a price of $95.70 per share. Assume that Chilly Moose Fruit new preferred shares can be sold without incurring flotation costs.

Chilly Moose Fruit does not have any retained earnings available to finance this project, so the firm will have to issue new common stock to help fund it. Its common stock is currently selling for $22.35 per share, and it is expected to pay a dividend of $2.78 at the end of next year. Flotation costs will represent 8% of the funds raised by issuing new common stock. The company is projected to grow at a constant rate of 8.7%, and they face a tax rate of 40%.

Chilly Moose Fruit’s WACC for this project will be: (Note: Round your answer to two decimal places.)

18.38%

15.32%

12.26%

14.55%
Answer #1
Peaceful Book Binding Company
Finding the WACC:
Before tax cost of Debt =9.6%
Tax rate =40%
After Tax cost of Debt =9.6%*(1-40%)=5.76%
WACC calculation
Type of Capital Market Value Weightage of Market Value Post Tax cost Weighted cost
Debt $        750,000 43.91% 5.76% 2.53%
Preferred Stock $          78,000 4.57% 10.70% 0.49%
Common Equity $        880,000 51.52% 13.50% 6.96%
Total $     1,708,000 100.00% 9.97%
PBBC WACC is 9.97%
Chilly Moose Fruit Producer
WACC Calculation
Cost of Debt
We need to find the YTM of bond which is the cost of it.
Face value of Bond $             1,000
Market Price of Bond $       1,555.38
Annual Coupn amount @11% $          110.00
Year to Maturity =n= 15
We need to Find the Yield to Maturity
YTM = [ Annual Interest +(Face Value-Market Price)/Years to maturity]/(Face value+2.Maketvalue)/3
YTM =[110+(1000-1555.38)/15]/(1000+2*1555.38)/3
YTM =5.4%
Tax rate =40%
After Tax cost of debt =5.4%*(1-40%)=3.24%
Cost of Preferred share
Estimated Price =$95.7
Annual Dividend =$8
Cost of preference share =8/95.7=8.36%
Common share cost
Current Price of common share = $22.35
Floatation cost per share @8% = $1.79
Net Realization per share = $20.56
Next year dividend =D1= $               2.78
Dividend growth rate =g= 8.70%
Assume cost of new equity =k
so P0=D1/(k-g)
20.56=2.78/(k-8.7%)
k=22.22%
So cost of Equity =22.22%
WACC Calculation
Type of Capital Weightage of Market Value Post Tax cost Weighted cost
Debt 35.00% 3.24% 1.13%
Preferred Stock 2.00% 8.36% 0.17%
Common Equity 63.00% 22.22% 14.00%
Total 100.00% 15.30%
So the closest answer is 15.32%
Chilly Moose Fruit's WACC is 15.32%